SINGLET-TRIPLET COUPLING IN ACETYLENE REVEALED IN DETAIL BY SURFACE ELECTRON EJECTION BY LASER EXCITED METASTABLES SPECTROSCOPY

\underline{RYAN L. THOM}, ADYA P. MISHRA, and R. W. FIELD, Department of Chemistry and G. R. Harrison Spectroscopy Laboratory, Massachusetts Institute of Technology, Cambridge, MA 02139.

We present here a new rotational analysis of the surface electron ejection by laser excited metastables (SEELEM) spectrum of acetylene in the energy region of the 3ν_{3} vibrational level of the Ã ^{1}A_{u} excited electronic state of C_{2}H_{2}. Experimental conditions have allowed for the first time the assignment of rotational quantum numbers to the long-lived, largely tripletcharacter SEELEM eigenstates, after which the doorway-mediated coupling of the triplet manifold to the singlet level can be fitted to a spin-orbit effective Hamiltonian. The results of this model concerning the vibronic identities of the triplet states, the magnitude of the spin-orbit interaction, and relative singlet/triplet SEELEM detectivities, will be discussed.